OpenStudy (anonymous):

Solve each quadratic equation. Show your work. x^2 + 3x = 10

OpenStudy (welshfella):

x^2 + 3x - 10 = 0 you need to find 2 numbers which when multiplied give -10 and when added give + 3

OpenStudy (welshfella):

then these go into 2 brackets as below (x + __)(x _ __)

OpenStudy (welshfella):

think about the factors of -10

OpenStudy (welshfella):

what comes to mind?

OpenStudy (anonymous):

2 and 5

OpenStudy (welshfella):

right - nearly there but its -10 not 10

OpenStudy (welshfella):

its either -2 and 5 or -5 and 2 right?

OpenStudy (welshfella):

which pair add up to +3 ?

OpenStudy (anonymous):

-2 and 5

OpenStudy (welshfella):

thats right so we write (x + 5)(x - 2) = 0 can you continue ?

OpenStudy (welshfella):

these are the factors of x^2 + 3x - 10 - now when 2 things are multiplied to give zero either one or the other must equal 0 so we have x + 5 =0 and x - 2 = 0 so we only have to solve these 2 small equations and we our answer these equations have 2 roots

OpenStudy (welshfella):

* quadratic equations have 2 roots

OpenStudy (anonymous):

so the answer is 2 or wat

OpenStudy (welshfella):

x - 2 = 0 so x = 2 x+5 = 0 so x = -5 the solution set is {-2,5}

OpenStudy (anonymous):

oh ok thanks alot

OpenStudy (welshfella):

yw